What is the force required for a steel wire to double it's length when it's area of cross section is one square cm. and Young's modules is 200Gpa.
1
Expert's answer
2021-10-18T10:59:40-0400

Solution.

"S=1cm^2=10^{-4}m^2;"

"E=200\sdot 10^9 Pa;"

"\Delta l=l_0;"

"F=k\Delta l;"

"k=E\sdot\dfrac{S}{l_0};"

"F=E\dfrac{S}{l_o}\sdot \Delta l=ES;"

"F=200\sdot 10^9\sdot 10^{-4}=200\sdot 10^5 N;"

Answer: "2\sdot10^7 N."